OpenStudy (anonymous):

If sec^2 A=5/2 What is 1-sin^2 A?

OpenStudy (anonymous):

sec =1/cos so cos=2/5

OpenStudy (anonymous):

so cos^2 a=2/5

OpenStudy (anonymous):

cos^2 a+ sin^2 a=1

OpenStudy (anonymous):

That was fast but didnt understand how you got it...

OpenStudy (anonymous):

cos^2a=1-sin^2a

OpenStudy (anonymous):

now got it

OpenStudy (anonymous):

transposed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

two identities were used here

OpenStudy (anonymous):

\[\sin ^2x+\cos^2x=1\]

OpenStudy (anonymous):

\[\sec^2=1/\cos^2a\]

OpenStudy (anonymous):

now u got it ?

OpenStudy (anonymous):

reciprocal and pythagorean identities....yup

OpenStudy (anonymous):

hmmmmm........

OpenStudy (anonymous):

dont know those identities

OpenStudy (anonymous):

?

Parth (parthkohli):

Do you understand this:\[1-\sin^2 A = \cos^2 A\]

OpenStudy (anonymous):

yes pythagorean

OpenStudy (anonymous):

so my answer here is just 2/5?

Parth (parthkohli):

Indeed.

OpenStudy (anonymous):

1-sin^2 A= 2/5

OpenStudy (anonymous):

?

Parth (parthkohli):

Indeed